A He–Ne laser illuminates a narrow, single slit that is 1850 nm wide. The first dark fringe is found at an angle of 20.0 degrees from the central peak.

1) Determine the wavelength of the light from the laser. (Express your answer to three significant figures.)
Answer #1

Let the distance from the center of central maxima be Y

Let the slit width be denoted by a

seperation between slit and screen be L

So,

y=Lλ/a

And we can say that,

y/Ltanθ=θ(for small θ)

So,

y/L=0.3488 radian...........(200=0.3488 radian)

So,

λ=aθ

=1850*0.3488

λ=645.444 nm